OpenStudy (anonymous):

if x^2+y^2 = z^2, show that xy/2 is not a perfect square. all are integers.

OpenStudy (dan815):

hmm

OpenStudy (dan815):

you have to show that x * y/2 cannot be the dimensions of a square if they satisfy this equation

OpenStudy (dan815):

in other words either x/2 = y or y/2=x

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

WLOG, we can say that x and y do not share a common factor \[2k^2=xy\] We also know that x|k and y|k So 2k^2/x is an [even] integer. Likewise 2k^2/y is an [even] integer. However, one of them must be odd.

OpenStudy (dan815):